# HD-GEN: A High-Performance Software System for Human Mobility Data Generation Based on Patterns of Life

### v1

@@ 4.1. Geolife+: Calibrated Dataset
To demonstrate the scalability of our framework, we generated multiple datasets based on the GeoLife trajectories, varying both the number of simulated agents and the simulation duration. These datasets range from small-scale runs with hundreds of agents to large-scale scenarios involving tens of thousands of agents, spanning months to several years. Compared to the original GeoLife dataset, our simulations capture a richer set of agent activities, as every agent is continuously observed throughout the simulation period. This produces datasets that are significantly larger in size while maintaining realistic mobility characteristics, offering researchers an expanded view of population-scale mobility (amiri2024geolife+).
@@ 4.2. Massive Trajectory Data
We generated a collection of large-scale check-in, location-based social network (LBSN) and trajectory datasets across four diverse urban and suburban regions: Fairfax (GMU campus area), New Orleans (French Quarter), San Francisco, and Atlanta. Each dataset integrates three components: simulated check-ins, evolving social links, and trajectory points recorded at five-minute intervals for every agent. Simulations were conducted at varying scales, from 1,000 to 10,000 agents over 15-month periods, as well as long-term runs spanning up to 20 years with 1,000 agents. These datasets capture both short-term dynamics and long-term behavioral patterns, providing a flexible foundation for mobility, social network, and anomaly detection research (amiri2023massive).
@@ 4.3. Human Mobility Dataset with Anomalies
To demonstrate the flexibility of HD-GEN, we generated datasets that incorporate controlled anomalies designed specifically for anomaly detection research. The simulation environment allows users to inject a wide range of behavioral deviations into agent activities, including irregular eating patterns, atypical social interactions, and disruptions to routine work behavior. Each anomaly can be configured at different intensity levels and applied either to individual agents or to the entire population, enabling the creation of diverse and realistic experimental scenarios. Agents can be designated as anomalous through three complementary mechanisms: (i) random assignment, in which anomalies are stochastically applied to selected agents; (ii) infectious disease style propagation, where anomalous behavior spreads dynamically through agent interactions, mimicking epidemiological processes; and (iii) location based triggers, where agents begin exhibiting anomalies after visiting specific locations. These mechanisms support the study of both isolated and systemic anomaly patterns and enable investigation into how abnormal behaviors emerge, propagate, and are detected in complex mobility systems (amiri2024urban).
@@
In related work, we introduced specific anomaly types to generate datasets for global anomaly detection tasks (zhang2024large; zhang2024transferable; liu2024neural). The implemented anomaly categories include hunger anomalies, social anomalies, and work anomalies. Hunger anomalies cause agents to experience hunger more frequently than normal, resulting in increased check ins at restaurants or homes. Social anomalies cause agents to ignore usual preferences and social influences, selecting recreational locations at random. Work anomalies disrupt routine behavior by causing agents to skip work on days when attendance is expected. Each anomaly type is further parameterized by three intensity levels: red, orange, and yellow. Red level anomalies represent extreme behavioral deviations, such as skipping work 100% of the time, orange level anomalies represent moderate deviation at approximately 50%, and yellow level anomalies represent mild deviation at approximately 20%. Anomalies can be applied either to specific agents for a defined duration or globally across all agents on selected days.
@@
Using the visualization tools described in 3.3, Fig. 11 presents an example of a global anomaly experiment. The figure shows check in patterns for multiple venue types, including apartments, pubs, restaurants, and workplaces, over a 1050 day simulation period starting in January 2024. The first 450 days simulate normal behavior, allowing agents to establish stable baseline routines. During the remaining 600 days, global anomalies are introduced, where all agents adopt a specific anomaly type and intensity level on each day. As illustrated in the figure, agent behavior remains stable during the baseline phase and becomes increasingly irregular once anomalies are introduced. For example, during red level global hunger anomaly events, all agents experience heightened hunger, leading to increased visits to restaurants and homes. This produces sharp spikes in restaurant and home check ins, accompanied by fluctuations in workplace visits as agents leave work to eat.
@@
We used the provided tools to generate the dataset reported in (amiri2024urban), which contains multiple anomaly types and configurations. The framework further supports flexible control over anomaly distribution by allowing users to select between random assignment, infectious propagation, and location based spreading mechanisms. Together, these capabilities enable the generation of rich, realistic datasets for evaluating anomaly detection methods across a wide range of behavioral and mobility scenarios.
@@ 5. Conclusions
In this paper, we have presented a software system to calibrate, generate, and process large-scale synthetic geospatial datasets by simulating the pattern of life of individuals with the consideration of real-world constraints. We have shown the effectiveness of our approach by generating a set of synthetic datasets based on the GeoLife dataset. Furthermore, we have shown that the generated datasets exhibit similar statistical properties to the original dataset, and can be used for various geospatial data analysis tasks. We have also provided detailed instructions on how to reproduce the generated datasets, and have made the code and data available on GitHub. With the provided datasets and code, researchers can easily generate large-scale synthetic geospatial datasets for their research purposes and evaluate the performance of their algorithms on realistic data. Our approach bridges the gap between real-world and simulated trajectory data by integrating statistical features from real datasets with the scalability of simulations. By introducing a data generation tool, a genetic algorithm for calibration, and a data processing pipeline, we enable the creation of realistic, large-scale mobility datasets suitable for diverse applications. This framework provides a robust foundation for research in human mobility modeling, machine learning, and data-driven applications while addressing the limitations of both real and synthetic datasets.

### v2

@@ 4.1. Geolife+: Calibrated Dataset
To demonstrate the scalability of our framework, we generated multiple datasets based on the GeoLife trajectories, varying both the number of simulated agents and the simulation duration. These datasets range from small-scale runs with hundreds of agents to large-scale scenarios involving tens of thousands of agents, spanning months to several years. Compared to the original GeoLife dataset, our simulations capture a richer set of agent activities, as every agent is continuously observed throughout the simulation period. This produces datasets that are significantly larger in size while maintaining realistic mobility characteristics, offering researchers an expanded view of population-scale mobility (Amiri et al., 2024c).
@@ 4.2. Massive Trajectory Data
We generated a collection of large-scale check-in, location-based social network (LBSN) and trajectory datasets across four diverse urban and suburban regions: Fairfax (George Mason University campus area), New Orleans (French Quarter), San Francisco, and Atlanta. Each dataset integrates three components: simulated check-ins, evolving social links, and trajectory points recorded at five-minute intervals for every agent. Simulations were conducted at varying scales, from 1,000 to 10,000 agents over 15-month periods, as well as long-term runs spanning up to 20 years with 1,000 agents. These datasets capture both short-term dynamics and long-term behavioral patterns, providing a flexible foundation for mobility, social network, and anomaly detection research (Amiri et al., 2023).
@@ 4.3. Human Mobility Dataset with Anomalies
To demonstrate the flexibility of HD-GEN, we generated datasets that incorporate controlled anomalies designed specifically for anomaly detection research. The simulation environment allows users to inject a wide range of behavioral deviations into agent activities, including irregular eating patterns, atypical social interactions, and disruptions to routine work behavior. Each anomaly can be configured at different intensity levels and applied either to individual agents or to the entire population, enabling the creation of diverse and realistic experimental scenarios. Agents can be designated as anomalous through three complementary mechanisms: (i) random assignment, in which anomalies are stochastically applied to selected agents; (ii) infectious disease style propagation, where anomalous behavior spreads dynamically through agent interactions, mimicking epidemiological processes; and (iii) location based triggers, where agents begin exhibiting anomalies after visiting specific locations. These mechanisms support the study of both isolated and systemic anomaly patterns and enable investigation into how abnormal behaviors emerge, propagate, and are detected in complex mobility systems (Amiri et al., 2024b).
@@
In related work, we introduced specific anomaly types to generate datasets for global anomaly detection tasks (Zhang et al., 2024a, b; Liu et al., 2024). The implemented anomaly categories include hunger anomalies, social anomalies, and work anomalies. Hunger anomalies cause agents to experience hunger more frequently than normal, resulting in increased check ins at restaurants or homes. Social anomalies cause agents to ignore usual preferences and social influences, selecting recreational locations at random. Work anomalies disrupt routine behavior by causing agents to skip work on days when attendance is expected. Each anomaly type is further parameterized by three intensity levels: red, orange, and yellow. Red level anomalies represent extreme behavioral deviations, such as skipping work 100% of the time, orange level anomalies represent moderate deviation at approximately 50%, and yellow level anomalies represent mild deviation at approximately 20%. Anomalies can be applied either to specific agents for a defined duration or globally across all agents on selected days.
@@
Using the visualization tools described in 3.4, Fig. 11 presents an example of a global anomaly experiment. The figure shows check in patterns for multiple venue types, including apartments, pubs, restaurants, and workplaces, over a 1050 day simulation period starting in January 2024. The first 450 days simulate normal behavior, allowing agents to establish stable baseline routines. During the remaining 600 days, global anomalies are introduced, where all agents adopt a specific anomaly type and intensity level on each day. As illustrated in the figure, agent behavior remains stable during the baseline phase and becomes increasingly irregular once anomalies are introduced. For example, during red level global hunger anomaly events, all agents experience heightened hunger, leading to increased visits to restaurants and homes. This produces sharp spikes in restaurant and home check ins, accompanied by fluctuations in workplace visits as agents leave work to eat.
@@
We used the provided tools to generate the dataset reported in (Amiri et al., 2024b), which contains multiple anomaly types and configurations. The framework further supports flexible control over anomaly distribution by allowing users to select between random assignment, infectious propagation, and location based spreading mechanisms. Together, these capabilities enable the generation of rich, realistic datasets for evaluating anomaly detection methods across a wide range of behavioral and mobility scenarios.
@@ 5. Conclusions
In this paper, we have presented a software system to generate, calibrate and process large-scale synthetic geospatial datasets by simulating the pattern of life of individuals with the consideration of real-world constraints. We have shown the effectiveness of our approach by generating a set of synthetic datasets based on the GeoLife dataset. Furthermore, we have shown that the generated datasets exhibit similar statistical properties to the original dataset, and can be used for various geospatial data analysis tasks. We have also provided detailed instructions on how to reproduce the generated datasets, and have made the code and data available on GitHub. With the provided datasets and code, researchers can easily generate large-scale synthetic geospatial datasets for their research purposes and evaluate the performance of their algorithms on realistic data. Our approach bridges the gap between real-world and simulated trajectory data by integrating statistical features from real datasets with the scalability of simulations. By introducing a data generation tool, a genetic algorithm for calibration, and a data processing pipeline, we enable the creation of realistic, large-scale mobility datasets suitable for diverse applications. This framework provides a robust foundation for research in human mobility modeling, machine learning, and data-driven applications while addressing the limitations of both real and synthetic datasets.